coin collection worth $2.90 contains nickels, dimes, and quaters. There are three times as many dimes as there are nickels, and the number of quarters is 2 more than the number of nickels
Answer by stanbon!(97)

Let number of nickels be "x".
Then number of dimes is "3x".
And number of quarters is "x+2".
Value of nickels is (5)x cents.
Value of dimes is (10)(3x)= 30x cents
Value of quarters is (25)(x+2)=25x+50 cents
EQUATION:
Value of nickels + Value of dimes + Value of quarters is 290 cents.
5x + 30x + 25x+50 = 290
60x=240
x= 4 (number os nickels)
3x= 12 (number of dimes)
x+2= 6 (number of quarters)